OpenStudy (specek18):

what's the numerator for the following rational expression? 5/q+9/q=?/q Will give medal! please help

OpenStudy (kropot72):

\[\large \frac{5}{q}+\frac{9}{q}=\frac{?}{q}\] You just need to add the two numerators on the left hand side of the equation to find the solution.

OpenStudy (specek18):

so the answer would be 14/q ?

OpenStudy (kropot72):

You are correct.